36) 1. An aqueous solution contains 0.219 M carbonic acid and 0.166 M nitric acid. Calculate...

36)

1.

An aqueous solution contains 0.219 M carbonic acid and 0.166 M nitric acid.

Calculate the carbonate ion concentration in this solution.
[CO32-] = ____mol/L.

2.


An aqueous solution contains 0.110 M carbonic acid and 0.146 M perchloric acid.

Calculate the carbonate ion concentration in this solution.
[CO32-] = _____mol/L.
Answer #1

1. Sol :-

ICE table of H2CO3 :

.........................H2CO3 (aq) ------------------------------> HCO3-(aq)..............+..................H+ (aq)

Initial (I).............0.219 M ..............................................0.0 M..........................................0.0 M

Change (C) .........- y ......................................................+ y ............................................+ y

Equilibrium (E).....(0.219-y) M............................................y M............................................y M

Expression of Ka1 is :

Ka1 = [H+].[HCO3-] / [H2CO3]

4.2 x 10-7 = y2/(0.219-y)

If y<<0.219, then neglect y as compare to 0.219.

4.2 x 10-7 x 0.219  = y2

y = 0.00030 M = [HCO3-]

ICE table of HCO3-:

.........................HCO3- (aq) ------------------------------> CO32-(aq)..............+..................H+ (aq)

Initial (I).............0.00030 M ..............................................0.0 M..........................................0.0 M

Change (C) .........- y ......................................................+ y ............................................+ y

Equilibrium (E).....(0.00030-y) M............................................y M............................................y M

Expression of Ka2 is :

Ka1 = [H+].[CO32-] / [HCO3-]

4.7 x 10-11 = y2/(0.00030-y)

If y<<0.219, then neglect y as compare to 0.00030 .

4.7 x 10-11 x 0.00030 = y2

y2 = 1.41 x 10-14

y = 1.19 x 10-7 M = [CO32-]

Now, ICF table when HNO3 reacts with CO32-

............................CO32- (aq)................+................HNO3 (aq) -----------------> HCO3- (aq)...........+............NO3- (aq)

Initial (I) ..............1.19 x 10-7 M .............................0.166 M .............................. 0.00030 M ...............................

Change (C)..........-1.19 x 10-7 M...........................- 1.19 x 10-7 M......................+1.19 x 10-7 M ..........................

Final (F)...................0.0 M......................................0.166 M (approx.).................0.00030 M(approx.).................

So,

[HCO3-] = 0.00030 M

and we already calculated the concentration of CO32- in 0.00030 M HCO3- = 1.19 x 10-7 M

Hence, Concentration of CO32- = [CO32-] = 1.19 x 10-7 M